A football is kicked vertically upward from the ground and a person gazing out of window sees it moving upwards past him at 5 m/s.Window is 15m above the ground.Ignore air resistance And take g=10.

How high does the ball go above ground? and how much time does it take to go from ground to highest point?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri Kartik

let the initial velocity of football is u m/sec

motion from ground to window

v=5m/sec

h=15m

v2 =u2 -2gh

25=u2 -2*10*15

u2=325

motion from window to highest point

u=5 m/sec , v=0

v2 =u2 -2gh

0=25-2*10*h

h=5/4 m

so total hight =15+5/4 =65/4 m

motion from ground to max hight

u=√325

v=0

v=u-gt

t=√325/10  sec


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin